OpenStudy (anonymous):

factor: ((1/x)-1) / (x-1)

OpenStudy (anonymous):

-1/x

OpenStudy (anonymous):

can you show me the steps in how you reached that answer?

OpenStudy (anonymous):

yes i can show u BOVICE

OpenStudy (anonymous):

LET US SOLVE FIRSTLY THE NUMERATOR: 1/x -1 = (1-x)/1 AND DENOMINATOR : (x-1) (1-x)/(x-1) that is -1/x

OpenStudy (anonymous):

have u got it

OpenStudy (anonymous):

IS IT CLEAR TO U IF YES THEN TELL ME

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

1/x -1 does not equal (1-x)/1

OpenStudy (anonymous):

why

OpenStudy (anonymous):

\[1 \div (x-1) \]

OpenStudy (anonymous):

no. reread the question. ((1/x)-1)/(x-1)

jhonyy9 (jhonyy9):

((1/x)-1)/(x-1)=((1-x)/x)/(x-1)=(1-x)/x x 1/(x-1) --- because (1/b)/c =(1/b)x(1/c)=1/bc --- hence you get (1-x)/x x 1/(x-1)=(-(x-1))/x x 1/(x-1) after you have simplified with (x-1) you get -1/x hope that you have understood it